OpenStudy (lisa123):

Find the center, vertices, and foci of the ellipse with equation x^2/400+ y^2 /256= 1

OpenStudy (lisa123):

@dan815 can you please help me?

OpenStudy (jdoe0001):

I assume you've covered ellipses already?

OpenStudy (lisa123):

yes

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{purple}{ a}}^2}+\cfrac{(y-{\color{blue}{ k}})^2}{{\color{purple}{ b}}^2}=1 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad vertices\ ({\color{brown}{ h}}\pm a, {\color{blue}{ k}}) \\ \quad \\ \cfrac{x^2}{400}+\cfrac{y^2}{256}=1\implies \cfrac{(x-0)^2}{20^2}+\cfrac{(y-0)^2}{16^2}=1\) see the center? notice, the bigger denominator is under the "x" that means the major axis is over the x-axis so is a horizontal ellipse so from the center, you move "a" units to the left and right and you'd find the vertices there

OpenStudy (lisa123):

Ok so is the center (-20,0) and (20,0)?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

so, to find the focus points, or foci well, again, we move from the center, over the x-axis but this time a difference distance we'll move to the left and right by "c" units and \(\bf c=\sqrt{a^2-b^2}\)

OpenStudy (lisa123):

c= sqrt 400-256 =144 144=12^2

OpenStudy (jdoe0001):

well... kinda... just 12 \(\bf c=\sqrt{400-256}\implies c=\sqrt{144}\implies c=12\)

OpenStudy (jdoe0001):



OpenStudy (lisa123):

OK great I understand! And how are you supposed to know if the major axis is for x or y if the problem doesn't say?

OpenStudy (jdoe0001):

for ellipses is defined by the "bigger denominator" 400 is bigger than 256 and "x" is above the 400 so the major axis runs over the x-axis for hyperbolas is defined by the "positive fraction" which is usually at the right-side anyway so if the positive fraction is the one with the "y" in it that means the hyperbola transversal axis runs over the y-axis if it's the "x" fraction then it runs over the x-axis

OpenStudy (jdoe0001):

hmm well. the positive fraction in hyperbolas, is usually at the left-side rather usually, doesn't have to, but usually

OpenStudy (lisa123):

Ok well thank you for all of your help! :)

OpenStudy (jdoe0001):

yw